OpenStudy (anonymous):

1. If tanx=−2.5 then tan(−x) = 2. If sinx=0.1 then sin(−x) = 3. If cosx=0.8 then cos(−x)= 4. If tanx=−0.5 then tan(pi+x)=

OpenStudy (anonymous):

or @hartnn

OpenStudy (anonymous):

whoever gets here first !!!

hartnn (hartnn):

\(\tan(-x)=-\tan x \\ \sin(-x)=-\sin x \\ \cos(-x)=\cos x\)

OpenStudy (anonymous):

why is cos +?

hartnn (hartnn):

because cos is even function, and for even functions, \(f(-x)=f(x)\)

hartnn (hartnn):

sin and tan are odd functions, where, \(f(-x)=-f(x)\)

OpenStudy (anonymous):

:O ahh yes i see... yes yes.

OpenStudy (anonymous):

continue wise one.

hartnn (hartnn):

\(\tan (x \pm n* \pi)=\tan x \\ so, \tan(x+\pi)= \tan x\)

OpenStudy (anonymous):

where the hell did that n come from? O_o

hartnn (hartnn):

that means for any value of n=1,2,3,4,.... i've first given generalized form, then have put n=1

OpenStudy (anonymous):

ahhh gotcha... you almost lost me for a sec.

hartnn (hartnn):

so you got all 4 now ?

OpenStudy (anonymous):

well, no because i am trying not to use my calculator. and honestly its a huge crutch for me in math so i seem rather lost without it... and so no.. no i dont plus i am kinda slow -_- and still dont know what to do.

hartnn (hartnn):

tanx=−2.5 tan (-x) = - tan x = - (-2.5) =... ?

OpenStudy (anonymous):

just 2.5?

hartnn (hartnn):

thats correct :) was it difficult ? try 2nd....

OpenStudy (anonymous):

:O ahhhhh soooo -.1, .8, ... hmm... pi-.5

OpenStudy (anonymous):

pi-.5 is incorrect... so .5 is 1/2... huh.. yeah idk what to do with that.

hartnn (hartnn):

tan (pi+x) = tan x = -0.5

OpenStudy (anonymous):

what happened to the pi?

OpenStudy (anonymous):

nevermind. it clicked ! thanks !

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

i will be back for you ! brb!!!